# AgenticTagger: Structured Item Representation for Recommendation with LLM Agents

### v1

@@ Abstract.
High-quality representations are a core requirement for effective recommendation. In this work, we study the problem of LLM-based descriptor generation, i.e., keyphrase-like natural language item representation generation frameworks with minimal constraints on downstream applications. We propose AgenticTagger, a framework that queries LLMs for representing items with sequences of text descriptors. However, open-ended generation provides little control over the generation space, leading to high cardinality, low-performance descriptors that renders downstream modeling challenging. To this end, AgenticTagger features two core stages: (1) a vocabulary building stage where a set of hierarchical, low-cardinality, and high-quality descriptors is identified, and (2) a vocabulary assignment stage where LLMs assign in-vocabulary descriptors to items. To effectively and efficiently ground vocabulary in the item corpus of interest, we design a multi-agent reflection mechanism where an architect LLM iteratively refines the vocabulary guided by parallelized feedback from annotator LLMs that validates the vocabulary against item data. Experiments on public and private data show AgenticTagger brings consistent improvements across diverse recommendation scenarios, including generative and term-based retrieval, ranking, and controllability-oriented, critique-based recommendation.
@@
Figure 1. Illustration of the proposed method. We represent items using an ordered sequence of coarse-to-fine LLM-generated natural language descriptors, directly extracting task-relevant information, such as genres for an item corpus of CDs and Vinyls.
@@ 1. Introduction
There have been numerous efforts in leveraging LLM for item representation in recommender systems. Typically, this process involves prompting LLMs for summarizations of item content (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c) or users’ behavioral sequence (Liu et al., 2025a, b; Jalan et al., 2024; Jiang et al., 2025; Wang et al., 2024d), which are then fed into a recommendation model as textual features. We study item-side, prompting-based discrete feature representation (i.e., natural language descriptor) generation with no constraints on the recommendation model architecture, as item representation is the foundation of effective recommendation (Rajput et al., 2023; Hou et al., 2023, 2025a, 2025b; Zhai et al., 2024; Cheng et al., 2023; Kang et al., 2021; Zhang et al., 2024b, 2025b). In particular, we seek frameworks capable of generating keyphrase-like descriptors, broadly applicable as ID or semantic features to a wide variety of downstream applications, complementary to existing efforts in building specialized recommendation frameworks exclusively for prompting LLMs for recommendations (Liu et al., 2025a, b; Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025).
@@
Figure 2. The AgenticTagger Framework. AgenticTagger exploits LLMs’ content-understanding ability to automatically construct an interpretable, hierarchical descriptor vocabulary from data. Then, relevant descriptors in the vocabulary could be assigned to corresponding items by querying LLMs in parallel over the item corpus. These features could then power various downstream applications, such as generative and term-based retrieval, ranking, and critique-based recommendation.
@@
We argue this setting holds three advantages complementary to its alternatives. Firstly, item-side representation scales more easily for the typical case where recommender systems serve more users than their items (McAuley et al., 2015; Gao et al., 2022). Secondly, prompt-based methods align well with modern language models’ language-based interaction paradigm (Gheorghe Comanici et, 2025; OpenAI, 2024), and scales well with improvements in base LLMs without the need for fine-tuning. Thirdly, text descriptors and sparse discrete features hold irreplaceable value in recommender system, due to their flexibility in benefiting diverse recommendation architectures (Zhang et al., 2024b, 2025b; Wang et al., 2017, 2021; Kabbur et al., 2013; Cheng et al., 2016; Song et al., 2019), as well as their efficiency and scalability when deployed in keyword-based retrieval settings (Wartena et al., 2010; Google Ads Help, 2026), which complements prior studies in adjacent directions such as LLM-augmented item embeddings (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025). Finally, as we discuss in later sections (Section 6.2), text-based descriptor representations naturally come with interpretability, and are well suited for user-facing applications.
@@
However, despite LLMs’ strong content understanding ability, LLM generations offer no control over its choice of vocabulary, while a good set of discrete features should ideally capture meaningful frequent patterns in the item corpus (Xie et al., 2025; Rajput et al., 2023; Singh et al., 2024). In particular, as recommender system corpora generally exceed the context limit of LLMs (McAuley et al., 2015; Harper and Konstan, 2015), it is challenging to incorporate the abstract constraint of generating descriptors that meaningfully highlights its similarities and differences w.r.t. other items, as this requires a global view of the entire item corpus. For example, we find many features produced by free-form generation only occur on a single item across the corpus (Section 6.1), whereas feature-based machine learning requires meaningfully frequent patterns to converge. As we later discuss in Section 7.3), this issue leads to vocabulary explosion, where large quantities of descriptors are produced, resulting in more unique features than items, making downstream feature modeling challenging, compromising recommendation quality.
@@
To this end, we propose AgenticTagger, a multi-LLM framework that queries LLMs for one-to-many natural language discrete features (i.e, tag-like descriptors, as shown in Figure 1) shared across items, where the descriptors come from an automatically learned low-cardinality vocabulary. Specifically, AgenticTagger features two key stages, (1) vocabulary building (Section 4.1), where AgenticTagger automatically learns a hierarchical coarse-to-fine descriptor taxonomy, drawing insights from LLM for taxonomy mining (Wan et al., 2024; Xie et al., 2025; Pham et al., 2024) and codebook learning for item representation in recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; He et al., 2025; Rajput et al., 2023); and (2) vocabulary assignment (Section 4.2), which prompts LLMs to assign relevant descriptors from the vocabulary to items, constraining their generation space. By pre-learning a set of high-quality, low-cardinality descriptors, AgenticTagger effectively constrains the LLMs’ generation space, equipping LLM with a global, hierarchical view of the item corpus, thereby improving generation quality.
@@
Under the general AgenticTagger framework, a core design challenge is how to effectively “propagate” signals through LLM generations during the vocabulary building stage, to effectively improve the quality of the vocabulary for covering diverse items in the item corpus. To this end, following the numerous recent successes in self-improving LLM agents through verbal feedback (Shinn et al., 2023; Majumder et al., 2024; Xie et al., 2024), we adopt an LLM-based refinement loop, which refines the vocabulary by aiming for better coverage of the item corpus through LLM-generated failure reports from items that do not fit into the current vocabulary. On the other hand, unlike these prior works, AgenticTagger uniquely adopts a multi-agent architect-annotator framework, where a stronger (yet slower) architect LLM maintains the descriptor vocabulary, while numerous copies of the weaker (yet faster) annotator LLMs scan through the item corpus in parallel, making AgenticTagger easily scalable to large datasets.
@@
We conduct experiments with AgenticTagger across public benchmarks and a private dataset, showing the broad applicability of AgenticTagger via its improvements to diverse recommendation settings including term-based retrieval, generative retrieval, and ranking with ID-like item features. We also show that these natural language descriptors natively bring interpretability and controllability to recommender systems, such as making generative recommender models compatible with critiques by controlling their decoding process via user feedback. Our contributions are as follows:
@@
  • We identify the core challenge in LLM-based item representation generation, namely, vocabulary mining, and propose a multi-agent framework for addressing such issue.
  • We show the value of LLM-agent-produced, ID-like natural language item representations in recommender systems, for their performance, interpretability, and flexibility.
@@ 2. Related Work
LLM-generated Features in RecSys. Prior work in introducing LLM-produced features in recommender system broadly falls into two streams: LLM-as-encoders (Zhang et al., 2024b, 2025b) and LLM-as-text-generators (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025; Liu et al., 2024, 2025a), and this work aims at enhancing the latter. In contrast to the popular goal of generating user profiles that improve an LLM itself at recommendation task (Liu et al., 2025a, b), we consider specifically the case where the recommendation model is a general feature-aware ML model, which aligns well with most existing deployed recommendation models. Compared to the most similar works that encode LLM-produced descriptions (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025), we are the first work to demonstrate the value of generating one-to-many discrete features, and the first work to demonstrate the value of LLM agents in feature generation for RecSys.
Discrete Representation in RecSys. Our work is also similar to a series of effort in building better discrete features for RecSys, which typically involve discretizing item representations with product quantization and its variants (Jegou et al., 2011; Ge et al., 2014; Hou et al., 2023, 2025a), RQ-VAE (van den Oord et al., 2017; Rajput et al., 2023), and k-means (Ju et al., 2025; Zhang et al., 2024c). However, these methods do not leverage the reasoning ability of LLMs, and as we will later show in the paper, discrete features produced from raw text by exploiting LLMs’ content understanding ability outperform their counterparts such as RQ-VAE-based semantic IDs.
LLM Agents for RecSys. There have been numerous recent efforts in leveraging LLM agents to improve RecSys. These frameworks generally focus on leveraging LLMs for item recommendation (Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025), or enabling better interactivity by incorporating LLM and LLM agents into the recommendation serving stage (Friedman et al., 2023; He et al., 2023; Zhang et al., 2024a). However, these frameworks typically require specialized pipelines, whereas AgenticTagger focuses on feature generation, which potentially benefits diverse downstream recommendation models. Meanwhile, to the best of our knowledge, AgenticTagger is the first framework to study agentic feature generation for recommender systems.
Table 1. Summary of Notations
Notation Description
$\mathcal{U},\mathcal{I}$ Sets of users and items
$S_{u}$ Interaction history for user $u\in\mathcal{U}$
$X_{i}$ Raw attributes for item $i\in\mathcal{I}$
$\Pi,\pi$ LLMs, and a specific instance
$\mathcal{E}$ Set of natural language error feedback
AA Descriptor feature vocabulary
$a$ A specific feature descriptor ($a\in A$)
$g(\cdot)$ Transformation mapping attributes to a subset of AA
$f_{\theta}$ Recommender system parameterized by $\theta$
@@ 3. Problem Statement
where $f_{\theta^{*}(g)}$ is the recommender model trained using the generated features, and $\mathcal{M}$ is a relevant performance metric, e.g., offline metrics and key business metrics. Note that while AgenticTagger do not directly optimize for end performance due to the challenge of propagating gradients over an API-supported LLM, the best vocabulary should ideally, in our problem formulation, maximize downstream task performance. We provide an overview of the notations used in Table 1.
@@ 4. AgenticTagger
As shown in Figure 2, the overall pipeline of AgenticTagger involves two key stages: a vocabulary building stage, where we produce a descriptor vocabulary by observing typical items in the dataset, and a vocabulary assignment stage that assigns the best set of descriptors to items.
@@ 4.1. Vocabulary Building
Output: Hierarchical feature set $A_{all}$
@@
$p_{0}\leftarrow\text{new Feature}(\text{items}\leftarrow\mathcal{I},\text{parent}\leftarrow\text{null})$
@@
During the vocabulary building stage, AgenticTagger automatically learns a vocabulary from the item corpus. However, gradient-based vocabulary learning (Rajput et al., 2023; Hou et al., 2023, 2025a; Zhang et al., 2024c) is inapplicable to prompting-based LLM agents, and thus common solution is prompting an LLM in mini-batches (Wan et al., 2024; Pham et al., 2024) and refining an in-context vocabulary via LLM self-refinement (Madaan et al., 2023). However, such an iterative approach involves a non-parallelizable loop over the whole item corpus, and scales poorly to large datasets (Xie et al., 2025).
@@
In light of the above issues, in AgenticTagger, we opt to scale existing LLM-based frameworks by making the vocabulary building process parallelizable, where multiple LLMs are deployed concurrently to update a shared vocabulary that describes observed items. An overview of this process is illustrated in Algorithm 1 and Algorithm 2. Specifically, the iterative feature generation follows a hierarchical-clustering-like pipeline (Algorithm 1), where we generate an initial set of features that partitions the whole item corpus, followed by a top-down procedure to generate finer-grained features that further partition items associated with each feature into smaller subsets. For example, if the initial features includes high-level category information such as “CDs and Vinyls”, all relevant items will then form a subcluster, for generating finer-grained features such as “Rock” and “Blues”, as shown in Figure 2-(i).
@@
At the core of this top-down generation loop is a multi-agent iterative refinement (Figure 2-(ii) and Algorithm 2) subroutine reused for each feature generation job $Q_{i}$, which aims to produce a set of features that describe common items in the corresponding subset. Our core insight is leveraging a stronger (though potentially slower) LLM, $\pi_{Architect}$, to generate and refine the vocabulary subset relevant to each $Q_{i}$, and copies of faster (and potentially weaker) LLMs, $\pi_{Annotator}$, which attempt to annotate the current item corpus using the available vocabulary and provide feedback to $\pi_{Architect}$ to improve the vocabulary. Compared to single-LLM self-refinement (Wan et al., 2024; Pham et al., 2024), the most costly $\pi_{Annotator}$ operations are parallelized, while we can still use their feedback to prompt $\pi_{Architect}$ to refine the vocabulary.
As shown in Algorithm 2, during the subroutine execution, we first prompt the $\pi_{\text{Architect}}$ LLM to generate the initial set of descriptors by observing a small subset of items $\text{Distill}(I_{sub})$ in its context, then deploy copies of (or batched queries to) the annotator LLMs $\pi_{\text{Annotator}}$ to determine if any of the currently known features apply to each item in $I_{sub}$, generating a set of error suggestions $\mathcal{E}$. The specific prompts are as shown in Section B.2. The error reports $\mathcal{E}$ then act inductive biases for prompting the vocabulary maintainer LLM $\pi_{\text{Architect}}$, to refine the current set of features. To avoid overcrowding $\pi_{\text{Architect}}$’s context, $\text{Distill}(\cdot)$ refers to a general operation that selects diverse representative samples from a larger set of item descriptions, such as K-Medoids clustering (Kaufman and Rousseeuw, 1987)’s centroids, as we use in our experiments. We repeat such an iterative refinement loop until a fixed number of optimization steps has been reached, or a desired number of items in $I_{sub}$ can be described by the generated features $A$. The union of features generated for each subset of the item corpus then forms the final feature space $A_{all}$, providing a coarse-to-fine description for the whole item corpus.

### v2

# AgenticTagger: Generating Structured Item Representation for Recommendation with LLM Agents

@@ Abstract.
High-quality representations are a core requirement for effective recommendation. In this work, we study the problem of LLM-based descriptor generation, i.e., keyphrase-like natural language item representation generation frameworks with minimal constraints on downstream applications. We propose AgenticTagger, a framework that queries LLMs for representing items with sequences of text descriptors. However, open-ended generation provides little control over the generation space, leading to high cardinality, low-performance descriptors that render downstream modeling challenging. To this end, AgenticTagger features two core stages: (1) a vocabulary-building stage in which a set of hierarchical, low-cardinality, and high-quality descriptors is identified, and (2) a vocabulary-assignment stage in which LLMs assign in-vocabulary descriptors to items. To effectively and efficiently ground vocabulary in the item corpus of interest, we design a multi-agent reflection mechanism in which an architect LLM iteratively refines the vocabulary guided by parallelized feedback from annotator LLMs that validate the vocabulary against item data. Experiments on public and private data show AgenticTagger brings consistent improvements across diverse recommendation scenarios, including generative and term-based retrieval, ranking, and controllability-oriented, critique-based recommendation.
@@
Figure 1. Illustration of the proposed method. Items are represented using an ordered sequence of LLM-generated descriptors that compress task-relevant information.
@@ 1. Introduction
High-quality item representations are the foundation of effective recommendation systems (Rendle et al., 2009; Kang et al., 2021; Rajput et al., 2023; Singh et al., 2024; Hou et al., 2023; Coleman et al., 2023; Zhang et al., 2024b, 2025b). These representations are commonly derived from collaborative filtering (Rendle et al., 2009; Koren et al., 2009; Kabbur et al., 2013a; He et al., 2017; Kang and McAuley, 2018) and/or content signals (Sharma et al., 2015; Mooney and Roy, 2000; Rajput et al., 2023; Singh et al., 2024; Hou et al., 2023, 2025b, 2025a; Wang et al., 2024e) via representation learning. Recently, there have been numerous attempts to enhance item representations with LLMs, typically by generating auxiliary unstructured item descriptions (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c) or prompting LLMs to infer user interests from interacted items and directly make recommendations (Liu et al., 2025a, b; Jalan et al., 2024; Jiang et al., 2025; Wang et al., 2024d). However, these approaches either require modeling high-cardinality unstructured natural language descriptions or deploying specialized recommendation frameworks, whereas large-scale recommender systems frequently benefit from low-cardinality, broadly applicable item representations such as hashed categorical features (Kang et al., 2021; Coleman et al., 2023; Singh et al., 2024; He et al., 2025; Hou et al., 2023, 2025b, 2025a). Thus, despite recent advances in both LLM-enhanced and structured item representations, it is unclear how to effectively leverage LLMs’ strong semantic understanding ability to enhance structured item representations.
@@
Figure 2. The AgenticTagger Framework. AgenticTagger exploits LLMs’ content-understanding ability to automatically construct an interpretable, hierarchical descriptor vocabulary from data. Then, relevant descriptors in the vocabulary are assigned to corresponding items by querying LLMs in parallel over the item corpus. These features could then power various downstream applications, such as generative retrieval, ranking, and critique-based recommendation.
@@
To this end, we study item-side, LLM-based generation frameworks for discrete feature representations (i.e., tag-like natural language descriptors), capable of exploiting LLMs’ reasoning ability to produce compact, structured, and fixed-slot nominal features with minimal constraints on downstream models, as shown in Figure 1. We argue that this setting holds three complementary values compared to its alternatives, such as generating unstructured item descriptions with LLMs or prompting LLMs directly for recommendations. First, discrete features such as hashed item IDs and text descriptors provide irreplaceable value in recommender systems, due to their compatibility with diverse architectures (Zhang et al., 2024b, 2025b; Wang et al., 2017, 2021; Kabbur et al., 2013b; Cheng et al., 2016; Song et al., 2019), their effectiveness in improving recommendation accuracy (Rajput et al., 2023; Hou et al., 2023, 2025b; Singh et al., 2024), and their scalability when deployed in keyword-based retrieval or advertisement targeting (Wartena et al., 2010; Google Ads Help, 2026). Second, item representation scales easily in the typical case where a recommender system serves more users than items, and the item corpus remains more stable than user interests (McAuley et al., 2015; Gao et al., 2022). Finally, as we discuss in a later section (Section 5.3), text-based descriptor representations provide inherent interpretability and are well-suited for user-facing applications.
@@
While existing efforts typically directly generate unstructured item descriptions from LLMs (Valizadeh et al., 2025; Ren et al., 2024; Shi et al., 2025; Yada and Yamana, 2024; Wang et al., 2024c), free-form LLM generations offer no control over their exact word choices, and thus there is no way to enforce structural constraints. For example, it is challenging to enforce mutual exclusivity between descriptions from free-form generation (Section 6.3), making these descriptors unsuitable for our use case. Further, a good set of text descriptors should ideally capture meaningful frequent patterns in the item corpus (Xie et al., 2025; Rajput et al., 2023; Singh et al., 2024). Yet, recommender system corpora generally exceed the context limit of LLMs (McAuley et al., 2015; Harper and Konstan, 2015), and thus it is challenging to incorporate the abstract constraint of generating descriptors that meaningfully highlight item similarities and differences w.r.t. other items, as this requires a global view of the entire item corpus. Finally, as LLMs can pick any arbitrary surface form (e.g., ”Blues music” and ”Blues great hits”) for the same underlying concept in free-form generation, such an approach often leads to vocabulary explosion, where most generated descriptors occur only once across the item corpus, as we discuss in Section B.3. This renders many features unusable without further processing, compromising downstream recommendation performance (Section 6.3).
@@
To address these issues, we propose AgenticTagger, a multi-LLM framework that queries LLMs for natural language descriptors (i.e., tag-like descriptors shown in Figure 1), where the descriptors come from an automatically learned low-cardinality, hierarchically structured vocabulary. Specifically, AgenticTagger features two key stages: (1) vocabulary building (Section 4.1), where AgenticTagger automatically learns a hierarchical coarse-to-fine descriptor taxonomy, drawing insights from LLMs for taxonomy mining (Wan et al., 2024; Xie et al., 2025; Pham et al., 2024) and codebook learning for item representation in recommendation (Hou et al., 2023, 2025b, 2025a; Singh et al., 2024; He et al., 2025; Rajput et al., 2023); and (2) vocabulary assignment (Section 4.2), which prompts LLMs to assign relevant descriptors from the vocabulary to items, constraining their generation space. By mining a high-quality, low-cardinality descriptor vocabulary and performing in-vocabulary generation, AgenticTagger effectively constrains the LLMs’ generation space while equipping LLMs with a global, hierarchical view of the item corpus, enforcing structural constraints on LLM-generated descriptors.
@@
Under the general AgenticTagger framework, a core design challenge is how to effectively “propagate” signals through LLM generations during the vocabulary-building stage to improve the quality of the vocabulary and better cover diverse items in the item corpus. To this end, following the numerous recent successes in self-improving LLM agents through verbal feedback (Shinn et al., 2023; Majumder et al., 2024; Xie et al., 2024), we adopt an LLM-based refinement loop, which refines the vocabulary by aiming for better coverage of the item corpus through LLM-generated failure reports on items that do not fit into the current vocabulary. Unlike these prior works, AgenticTagger uniquely adopts a multi-agent architect-annotator framework, where a stronger (yet slower) architect LLM maintains the descriptor vocabulary, while numerous copies of the weaker (yet faster) annotator LLM scan through the item corpus in parallel, making AgenticTagger scalable.
@@
We conduct experiments on AgenticTagger across public benchmarks and a private dataset, showing the broad applicability of AgenticTagger via its improvements to diverse recommendation settings, including generative recommendation and ranking. We also show that these natural language descriptors natively bring interpretability and controllability to recommender systems, such as by making generative recommender models compatible with critiques by controlling their decoding process via user feedback. Our contributions are as follows:
@@
  • We identify a core challenge in LLM-based structured item representation generation, namely, the lack of vocabulary control, and propose a multi-agent, two-stage descriptor generation framework to address this issue.
  • We show the value of LLM-agent-produced, structured natural language item representations in recommender systems, for their performance, interpretability, and flexibility.
@@ 2. Related Work
LLM-generated Features in RecSys. Prior work in introducing LLM-produced features in recommender systems broadly falls into two streams: LLM-as-encoders (Zhang et al., 2024b, 2025b) and LLM-as-text-generators (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025; Liu et al., 2024, 2025a), and this work aims at enhancing the latter. In contrast to the popular goal of generating user profiles that improve an LLM itself at the recommendation task (Liu et al., 2025a, b), we consider specifically the case where the recommendation model is a general feature-aware ML model, which aligns well with most existing deployed recommendation models. Compared to the most similar works that encode LLM-produced descriptions (Ren et al., 2024; Yada and Yamana, 2024; Valizadeh et al., 2025), we are the first to study fixed-slot, nominal feature generation, and the first to demonstrate the value of LLM agents in feature generation for RecSys.
Discrete Representation in RecSys. Our work is also similar to a series of efforts in building better discrete features for RecSys, which typically involve discretizing item representations with product quantization and its variants (Jegou et al., 2011; Ge et al., 2014; Hou et al., 2023, 2025a), RQ-VAE (van den Oord et al., 2017; Rajput et al., 2023), and K-means (Ju et al., 2025; Zhang et al., 2024c). However, these methods do not leverage the reasoning ability of LLMs. As we will later show in the paper, discrete features produced from raw text by exploiting LLMs’ content understanding ability outperform their counterparts, such as RQ-VAE-based semantic IDs.
LLM Agents for RecSys. There have been numerous recent efforts in leveraging LLM agents to improve RecSys. These frameworks generally focus on leveraging LLMs for item recommendation (Shu et al., 2023; Zhao et al., 2024; Thakkar and Yadav, 2024; Wang et al., 2024b; Xia et al., 2026; Huang et al., 2025), or enabling better interactivity by incorporating LLMs and LLM agents into the recommendation-serving stage (Friedman et al., 2023; He et al., 2023; Zhang et al., 2024a). However, these frameworks typically require specialized pipelines, whereas AgenticTagger focuses on feature generation, which potentially benefits diverse downstream recommendation models. Meanwhile, to the best of our knowledge, AgenticTagger is the first framework to study agentic feature generation for recommender systems.
@@ 3. Problem Statement
where $f_{\theta^{*}(g)}$ is the recommender model trained using the generated features, and $\mathcal{M}$ represents relevant performance metrics, e.g., offline metrics and key business metrics. Note that while AgenticTagger does not directly optimize for end performance due to the challenge of propagating gradients through an API-supported LLM, the best vocabulary should ideally, in our problem formulation, maximize downstream task performance. We provide an overview of the notations used in Table 10.
@@ 4. AgenticTagger
As shown in Figure 2, the overall pipeline of AgenticTagger involves two key stages: the vocabulary-building stage, where we produce a descriptor vocabulary by observing typical items in the dataset, and the vocabulary-assignment stage, where we assign the best set of descriptors to each item.
@@ 4.1. Vocabulary Building
Output: Hierarchical descriptor set $A_{all}$
@@
$p_{0}\leftarrow\text{new Descriptor}(\text{items}\leftarrow\mathcal{I},\text{parent}\leftarrow\text{null})$
@@
During the vocabulary-building stage, AgenticTagger automatically learns a vocabulary from the item corpus. However, gradient-based vocabulary learning (Rajput et al., 2023; Hou et al., 2023, 2025a; Zhang et al., 2024c) is inapplicable to prompting-based LLM agents, and thus the common solution is prompting an LLM in mini-batches (Wan et al., 2024; Pham et al., 2024) and refining an in-context vocabulary via LLM self-refinement (Madaan et al., 2023). Yet, such an iterative approach involves a non-parallelizable loop over the whole item corpus, as LLMs must sequentially refine their in-context vocabulary after seeing each batch of items, which scales poorly to large datasets (Xie et al., 2025).
@@
In light of the above issues, in AgenticTagger, we opt to scale existing LLM-based frameworks by making the vocabulary-building process parallelizable, where multiple LLMs are deployed concurrently to update a shared vocabulary that describes observed items. An overview of this process is illustrated in Algorithm 1 and Algorithm 2. Specifically, the iterative feature generation follows a hierarchical-clustering-like pipeline (Algorithm 1), where we generate an initial set of features that partitions the whole item corpus, followed by a top-down procedure to generate finer-grained features that further partition items associated with each feature into smaller subsets. For example, if the initial features include high-level category information such as “CDs and Vinyl”, all relevant items will then form a subcluster, for generating finer-grained features such as “Rock” and “Blues”, as shown in Figure 2-(i). By decomposing the task of generating an entire vocabulary into a top-down hierarchical procedure, AgenticTagger effectively limits the number of descriptors an LLM needs to produce at a time, which scales better to a large corpus. Meanwhile, the resulting hierarchical vocabulary naturally organizes items in a coarse-to-fine taxonomy, similar to hierarchical semantic IDs (Rajput et al., 2023; He et al., 2025), which are known to act as effective item representations.
@@
At the core of this top-down generation loop is a multi-agent iterative refinement (Figure 2-(ii) and Algorithm 2) subroutine reused for each feature generation job $Q_{i}$, which aims to produce a set of features that describe common items in the corresponding subset. Our core insight is leveraging a stronger (though potentially slower) LLM, $\pi_{Architect}$, to generate and refine the vocabulary subset relevant to each $Q_{i}$, and copies of a faster (and potentially weaker) LLM, $\pi_{Annotator}$, which attempt to annotate the current item corpus using the available vocabulary and provide feedback to $\pi_{Architect}$ to improve the vocabulary. Unlike in single-LLM self-refinement (Wan et al., 2024; Pham et al., 2024; Madaan et al., 2023), the most costly $\pi_{Annotator}$ operations are parallelized in AgenticTagger, while we can still use their feedback to prompt $\pi_{Architect}$ to refine the vocabulary.
As shown in Algorithm 2, during the subroutine execution, we first prompt the $\pi_{\text{Architect}}$ LLM to generate the initial set of descriptors by observing a small subset of items, $\text{Distill}(I_{sub})$, in its context, and then deploy copies of (or batched queries to) the annotator LLMs $\pi_{\text{Annotator}}$ to determine if any of the currently known features apply to each item in $I_{sub}$, generating a set of error suggestions $\mathcal{E}$. The specific prompts are as shown in Section B.1. The error reports $\mathcal{E}$ then act as inductive biases for prompting the vocabulary maintainer LLM $\pi_{\text{Architect}}$ to refine the current set of features. To avoid overcrowding $\pi_{\text{Architect}}$’s context, $\text{Distill}(\cdot)$ refers to a general operation that selects diverse representative samples from a larger set of item descriptions, such as the centroids from K-Medoids clustering (Kaufman and Rousseeuw, 1987), as we use in our experiments. We repeat this iterative refinement loop until a fixed number of optimization steps has been reached, or a desired number of items in $I_{sub}$ can be described by the generated features $A$. The union of features generated for each subset of the item corpus then forms the final feature space $A_{all}$, providing a coarse-to-fine description for the whole item corpus.